OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)? negative three halves negative two thirds three halves two thirds

OpenStudy (anonymous):

rise over run; y=mx+b

OpenStudy (anonymous):

so

OpenStudy (anonymous):

(x1-x2)/(y1-y2) x1=-2; x2=2; y1=4; y2=-2

OpenStudy (anonymous):

Let me know if you need it done further.

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

What is the x value of the first ordered pair?

Directrix (directrix):

This slope formula is not correct. >> (x1-x2)/(y1-y2) @rosho

OpenStudy (anonymous):

Sorry, been awhile since I've done algebra.

Directrix (directrix):

Denote (-2, 4) as (x1, y1) and (2, -2) as (x2, y2). The numerator of the slope is --> (-2) - 4 = -2 - 4 = -6. The denominator of the slope is --> (-2) - (-2) = -2 + 4 = 2. slope = m = (-6) / 2 = -3/2.